Find the equation of the clastic curve and its maximum deflection for the simply supported beam of length 27, having uniformly distributed load W per unit length.


1
Expert's answer
2022-01-05T05:26:02-0500

The maximum deflection of a simply supported beam at midspan is given by;

"\delta=\frac{5WL^4}{384EI}"

"\delta=\frac{5W27^4}{384EI}"

"=\frac{6919.80W}{EI}"